OpenStudy (anonymous):

Mrs. Computer is 3 times older than her daughter, Mousy. The sum of their ages is 52. How old is Mousy?

OpenStudy (diyadiya):

Let mousy's age be x What's Mrs .Computers age ?

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

3x+x=52 mrs computer is 39. her daughter is 13

OpenStudy (diyadiya):

3 times older means \( 3 \times x \) so 3x

OpenStudy (anonymous):

yes 4x=52 then divide both side by 4 and you get x then plug x into the initial equation 3x+x=52

OpenStudy (anonymous):

ok i get it now i just need help to put it in a equation

OpenStudy (anonymous):

okay so x=13 because 52/4=13 so it should look like this 3(13)+13=52 39+13=52

OpenStudy (anonymous):

therefore the age of the mother is 39 and the age of the daughter is 13

OpenStudy (anonymous):

thank u

OpenStudy (anonymous):

your welcome